### v1

# Exploring the Non-Markovian Dynamics in Depolarizing Maps

Abstract
The non-Markovian depolarizing channel is explored from the perspective of understanding its non-Markovian behaviour as well as the occurrence of singularities. The study brings together the various ways to identify and quantify non-Markovianity. This includes dynamical techniques such as quantum information backflow witness, BLP, RHP and HCLA measures. In addition, geometrical visualization of non-Markovian effects is presented using the variation in the volume of accessible states during dynamical evolution. Further, a trajectory-based visualization of the dynamical map, of the depolarizing channel, within the parameter space is presented. The trajectories traced during evolution, demonstrate the loss of CP-divisibility and the emergence of non-Markovianity under systematic variations of the system parameters.
preprint: APS/123-QED
I Introduction
The dynamical evolution of the open quantum system [1, 2], a composite system consisting of a smaller sub-system of interest and an interacting larger sub-system called environment, is of interest to the field of quantum information [3, 4] due to arising memory effects and non-Markovianity [5, 6, 7, 8, 9, 10, 11, 12]. The concept of the completely positive trace preserving (CPTP) map is used to study such evolution, which maps the system’s initial state $\rho_{S}(0)$ to the evolved final state $\rho_{S}(t)$ [3, 13]
$\rho_{S}(t)=\Phi(t,0)\rho_{S}(0).$ (1)
Alternately, CPTP maps are called quantum channels and can be described using operator sum representation (Kraus formalism) [14, 15, 16, 17]. The mapped evolution by the CPTP map can be classified as Markovian or non-Markovian. The Markovian evolution implies that system evolution is independent of past events and is dependent on only the present. On the contrary, such independence is violated, or the environment is found to pertain to the history of system evolution for non-Markovian evolution.
To identify non-Markovianity, traditionally, one resorts to the loss of CP-divisibility [7] or non-monotonic increase in the distinguishability of a pair of states during evolution [6]. Even CP-divisible processes can exhibit non-Markovain effects [9, 10]. The increase in distinguishability is indicative of information backflow from environment to system, indicating non-Markovianity. The Breuer-Laine-Piilo (BLP) measure [5] of non-Markovianity uses distinguishability to quantify non-Markovianity. CP-divisibility allows dynamics to be studied in discrete time intervals; the inability to do so is characteristic of a CP-indivisible or non-Markovan process. The dynamical map is said to be CP-indivisible, if the intermediate map, mapping intermediate evolution, is not completely positive (NCP), possessing at least one negative eigenvalue for the corresponding Choi matrix. The intermediate NCP map is indicative of system-environment correlation. The Rivas-Hulga-Plenio (RHP) measure [18, 7] uses the previously mentioned criteria to quantify non-Markovianity. Similarly, the Hall-Creser-Li-Anderson (HCLA) measure [19] quantifies non-Markovian effects but relies on negative time-dependent decay rates, which are indicative of non-Markovianity. Both measures are equivalent to each other and provide similar insights into non-Markovian processes.
Here, we undertake a follow-up of [20] to further explore the non-Markovian depolarising map. The additional emphasis on the quantum nature of information backflow [21] and originating singularities is presented. The singularity structure originates in the intermediate map at the crossover of two eigenvalues and is one of the features of depolarizing maps.
Our goal in this article is to answer the following two questions: (1) Under what conditions the depolarizing map becomes non-Markovian?, (2) Can we quantify the amount of this non-Markovianity?
In order to answer these questions, both dynamical, based on CP-divisibility, distinguishability, negative time-dependent decay rates, quantum information backflow witness, and geometrical techniques are employed, and a comprehensive understanding of the non-Markovianity of depolarizing maps is attempted. The geometrical visualisation involves descriptions of the non-Markovianity based on: an unexpected increase in the volume of accessible states of the system signifying the non-Markovian dynamics [22], and employing a trajectory-based visualisation of traced trajectories within the parameter space to capture the dynamical evolution of the depolarizing channel, in turn demonstrating the loss of CP-divisibility and the emergence of non-Markovianity under systematic variations in the system parameters [23].
This work is organized as follows. In Section II, the general depolarizing map including the non-Markovianity parameters is derived. In section III, we describe the dynamical techniques in detecting the non-Markovianity of the depolarizing map by observing the negativity of the eigenvalues of the Choi matrix of the intermediate map, witness for quantum information backflow, and the canonical time-dependent decay rate. Furthermore, we utilize the trace-distance-based distinguishability method to detect the non-Markovianity of the map. In section IV, we describe the geometrical tools in detecting the non-Markovianity through: (a) changing the volume of physical states of the system, (b) employing a trajectory-based visualization within the parameter space to trace trajectories and capture the dynamical evolution of the depolarizing channel. Finally, we conclude in section V.
II General depolarizing map
The depolarizing map is frequently encountered in quantum information theory. It takes an initial state of the system $\rho_{S}(0)$ to a mixture of itself and a maximally mixed state [3, 24, 25, 26, 20]. Mathematically, it can be represented as a quantum channel or quantum operation. In general, the operator sum representation of this map is $\Phi(\rho)=\sum_{i}E_{i}\rho E_{i}^{\dagger}$, where the Kraus operators are:
$E_{\mathbb{I}}=\sqrt{1-\frac{3}{4}k}\hskip 5.69054pt\mathbb{I};\hskip 19.91684ptE_{X}=\sqrt{\frac{k}{4}}\hskip 5.69054pt\sigma_{X},$
[…]
where $\mathbb{I}$ is the identity operator, and $\sigma_{X}$, $\sigma_{Y}$, $\sigma_{Z}$ are the Pauli matrices. Furthermore, $k$ increases monotonically from 00 (noiseless case) to $1$ (maximal depolarizing) [27]. The most general form of the Kraus operators of the depolarizing map can be written as [28, 20, 27]
$E_{\mathbb{I}}$ $=\sqrt{\left[1+\Upsilon_{1}(p)\right]\left(1-\frac{3}{4}p\right)}\hskip 5.69054pt\mathbb{I};$
$E_{X}$ $=\sqrt{\left[1+\Upsilon_{2}(p)\right]\frac{p}{4}}\hskip 5.69054pt\sigma_{X};$
$E_{Y}$ $=\sqrt{\left[1+\Upsilon_{2}(p)\right]\frac{p}{4}}\hskip 5.69054pt\sigma_{Y};$
$E_{Z}$ $=\sqrt{\left[1+\Upsilon_{2}(p)\right]\frac{p}{4}}\hskip 5.69054pt\sigma_{Z},$ (3)
where $\Upsilon_{1}(p)$ and $\Upsilon_{2}(p)$ are real functions, and $p$ is a time-like parameter that changes monotonically from 00 to $1$. We regain Eq. (II) by adjusting $\Upsilon_{1}(p)=\Upsilon_{2}(p)=0$, and $p$ being replaced by $k$.
To find out the form of $\Upsilon_{1}(p)$ and $\Upsilon_{2}(p)$, we use the completeness condition
$\sum_{i}E^{\dagger}_{i}E_{i}$ $=\left(1+\Upsilon_{1}(p)\right)\left(1-\frac{3}{4}p\right)+\left(1+\Upsilon_{2}(p)\right)\left(\frac{3}{4}\right)p=1,$
$\implies\left(1-\frac{3}{4}p\right)\Upsilon_{1}(p)+\frac{3}{4}p\Upsilon_{2}(p)=0.$ (4)
This suggests that $\Upsilon_{1}(p)=-\frac{3}{4}\alpha p$ and $\Upsilon_{2}(p)=\alpha\left(1-\frac{3}{4}p\right)$, where $\alpha$ is real number. The parameter $\alpha$ creates a small perturbation to the map and imposes non-Markovianity into the dynamics. Hence, it can be considered as a non-Markovian parameter.
The general Kraus operators of the depolarizing map then take the form
$E_{\mathbb{I}}$ $=\sqrt{\left[1-\frac{3}{4}\alpha p\right]\left(1-\frac{3}{4}p\right)}\hskip 5.69054pt\mathbb{I};$
$E_{X}$ $=\sqrt{\left[1+\alpha\left(1-\frac{3}{4}p\right)\right]\frac{p}{4}}\hskip 5.69054pt\sigma_{X};$
$E_{Y}$ $=\sqrt{\left[1+\alpha\left(1-\frac{3}{4}p\right)\right]\frac{p}{4}}\hskip 5.69054pt\sigma_{Y};$
$E_{Z}$ $=\sqrt{\left[1+\alpha\left(1-\frac{3}{4}p\right)\right]\frac{p}{4}}\hskip 5.69054pt\sigma_{Z}.$ (5)
To validate the map’s complete positivity, we choose $\alpha\in[0,1]$. For instance, if $\alpha=0$, then Eq. (II) reduces to the Kraus operators of the standard depolarizing map, Eq. (II), and $p$ becomes $k$.
In general, $k$ depends on $p$. Comparing Eq. (II) with Eq. (II), $k(p)$ can be expressed as
$k(p)=p+\alpha p-\frac{3}{4}\alpha p^{2}.$ (6)
III Non-Markovianity using dynamical tools
In this section we characterize the non-Markovianity of the depolarizing map, Eq. (II), by observing the negative eigenvalues of the Choi matrix of the intermediate map and the canonical time-dependent decay rate. In addition to it, the witness operator for non-Markovianity is discussed, and the quantification of resulting memory effects due to information backflow is also presented. Furthermore, we utilize the trace-distance-based distinguishability method to detect this non-Markovianity.
If the dynamical map $\Phi(t,0)$ in Eq. (1), is CP-divisible and also invertible, $i.e.,$ the inverse $\Phi^{-1}(t,0)$ exists for all times from $t=0$ to $t$, then one can write the dynamical map for $t\geq s\geq 0$ into the following composition [7],
$\Phi(t,0)=\Phi(t,s)\circ\Phi(s,0),$ (7)
where the $\circ$” represents the composition between two maps. The map $\Phi(t,s)$ describes the intermediate system-environment interactions and can be defined as
$\Phi(t,s)=\Phi(t,0)\left[\Phi^{-1}(s,0)\right],$ (8)
and needs to be CP. Note that the condition in Eq. (7) is the quantum counterpart to the classic Chapman-Kolgomorov equation [7]. However, if the intermediate map $\Phi(t,s)$ is not CP (NCP) [29], then it is indicative of the usual notion of non-Markovianity [5, 7, 20].
The dynamical map $\Phi(t,0)$ in terms of “divisibility”, represents a memoryless evolution, as a composition of physical maps is indicative of quantum Markovianity. Based on the RHP criterion [7], a dynamics is said to be non-Markovian if it is not CP divisible. This does not require optimization but a normalization is needed to handle the singularity. An alternative method of characterizing non-Markovian dynamics is the BLP method [5, 6]. Here, non-Markovianity is defined in terms of the non monotonic behavior of the trace distance, ||.$||.||_{1},$ and is based on the notion of distinguishability between any two initial states $\rho_{1}$ and $\rho_{2}$. A non monotonic behavior of distinguishability is indicative of a back-flow of information from the environment to the system, a signature of non-Markovianity. This requires optimization of the states and does not require normalization. The information backflow criteria does not distinguish between classical and quantum memory effects. Using [21], quantumness in the observed information backflow can be characterised for the intermediate map.
III.1 The eigenvalues of the Choi matrix of the intermediate dynamics
In general, the dynamical map $\Phi(t,0),t>0$ can be obtained by tomography [3, 7]. For the present case, the general depolarizing map is theoretically known, see Eq. (II). To realize the non-Markovianity of the general depolarizing map Eq. (II), we consider the intermediate map $\Phi(t,s)$, $0\leq s\leq t$. If the intermediate map $\Phi(t,s)$ is not CP, the dynamics will be non-Markovian [7].
In the following, we describe the main steps toward calculating the intermediate map in Eq. (8), $\Phi(t,s)$. The general depolarizing map is written as
$\Phi(t,0)(\rho(0))$ $=E_{\mathbb{I}}\rho(0)E^{\dagger}_{\mathbb{I}}+\sum_{i=X,Y,Z}E_{i}\rho(0)E_{i}^{\dagger}$
$=\left[1-\frac{3}{4}\alpha p\right]\left(1-\frac{3}{4}p\right)\hskip 2.84526pt\rho(0)+$
$\sum_{i=X,Y,Z}\left[1+\alpha\left(1-\frac{3}{4}p\right)\right]\frac{p}{4}\hskip 2.84526pt\sigma_{i}\rho(0)\sigma_{i}.$ (9)
Furthermore, we vectorize $\Phi(t,0)$ by writing the resulting matrix as a (column) vector by gathering the columns on top of one another. This process is called “vectorization” [7, 30, 31]. The diagonalized form of the matrix elements of the unital intermediate map obtained using $\Phi(t,0)$, following Eq. (8) is
$\Phi(t,s)=\Phi(t,0)\Phi^{-1}(s,0)=\begin{pmatrix}1&0&0&0\\ 0&\lambda_{1}(p)&0&0\\ 0&0&\lambda_{2}(p)&0\\ 0&0&0&\lambda_{3}(p)\end{pmatrix},$ (10)
where,
$\lambda_{1}(p)=\lambda_{2}(p)=\lambda_{3}(p)=\frac{p\left(4+4\alpha-3\alpha p\right)-4}{4q+4\alpha q-3\alpha q^{2}-4}.$ (11)
Here, $0\leq q\leq p\leq 1$. It is worth mentioning here that when $q=\alpha=0$ in Eq. (11), then $\Phi(t,s)=\Phi(t,0)$, and $\lambda_{1}(p)=\lambda_{2}(p)=\lambda_{3}=(1-p)$. This illustrates the standard depolarizing map, and the map shrinks the Bloch sphere uniformly along $x$, $y$, and $z$ to a radius of $(1-p)$ [26, 24].
To find out whether $\Phi(t,s)$ is NCP or not, we use a Choi matrix constructed by first constructing the matrix $U_{2\leftrightarrows 3}[\Phi(t,s)\otimes\mathbb{I}_{4}]U_{2\leftrightarrows 3}$ where $U_{2\leftrightarrows 3}$ is the commutation (or ”swap“) matrix between the ”second“ and the ”third“ subspace [111]; second, apply $U_{2\leftrightarrows 3}[\Phi(t,s)\otimes\mathbb{I}_{4}]U_{2\leftrightarrows 3}$ on $vec(\ket{\Psi}\bra{\Psi})$, where |$\ket{\Psi}=\frac{1}{\sqrt{2}}\left(\ket{00}+\ket{11}\right)$. $U_{2\leftrightarrows 3}$ is written as follows
$U_{2\leftrightarrows 3}=\mathbb{I}_{2}\otimes\begin{pmatrix}1&0&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{pmatrix}\otimes\mathbb{I}_{2}.$ (12)
We assume that $\phi(t,s)\otimes\mathbb{I}_{4}$ acts in a tensor product of four spaces with the same dimension , $\mathcal{H}_{1}\otimes\mathcal{H}_{2}\otimes\mathcal{H}_{3}\otimes\mathcal{H}_{4}$. Then $U_{2\leftrightarrows 3}$ denotes the permutation matrix interchanging the second and third subspace, $i.e.$ $U_{2\leftrightarrows 3}=\mathbb{I}_{2}\otimes U_{P}\otimes\mathbb{I}_{2}$. Third, we write the result as a matrix, $i.e.$ “devectorize” to construct Choi matrix of the intermediate map as follows
$\chi(\alpha,q,p)=\left[\Phi(t,s)\otimes\mathbb{I}\right]\ket{\Psi}\bra{\Psi}.$ | . (13)
By Choi-Jamiolkowski isomorphism, matrix $\chi(\alpha,q,p)$ is positive if $\Phi(t,s)$ is CP [20, 32]. An NCP $\Phi(t,s)$ is indicative of non-Markovian dynamics. The Choi matrix for the intermediate map, $\chi(\alpha,q,p)$, is found to be
$\chi(\alpha,q,p)=\begin{pmatrix}\frac{p+q+\alpha(p+q)-\frac{3}{4}\alpha(q^{2}+p^{2})-2}{q+\alpha q-\frac{3}{4}\alpha q^{2}-1}&0&0&\frac{(1+\alpha)p-\frac{3}{4}\alpha p^{2}-1}{2(1+\alpha)q-\frac{3}{2}\alpha q^{2}-4}\\ 0&\frac{(q-p)\left(\alpha-\frac{3}{4}(p+q)+1\right)}{q\left(1+\alpha\right)-\frac{3}{4}\alpha q^{2}-1}&0&0\\ 0&0&\frac{(q-p)\left(\alpha-\frac{3}{4}(p+q)+1\right)}{q\left(1+\alpha\right)-\frac{3}{4}\alpha q^{2}-1}&0\\ \frac{(1+\alpha)p-\frac{3}{4}\alpha p^{2}-1}{2(1+\alpha)q-\frac{3}{2}\alpha q^{2}-4}&0&0&\frac{p+q+\alpha(p+q)-\frac{3}{4}\alpha(q^{2}+p^{2})-2}{q+\alpha q-\frac{3}{4}\alpha q^{2}-1}\end{pmatrix}.$ (14)
Figure 1: (Color online). The eigenvalues of the Choi matrix (14). Solid, black line($\lambda_{\mathbb{I}}$) and, dashed, black line, ($\lambda_{X,Y,Z}$), both for $\alpha=0.0$. Solid, red line ($\lambda_{\mathbb{I}}$) and, red, dashed line ($\lambda_{\mathbb{I}}$) both for $\alpha=0.7$. The intermediate $p$-range lies between $p=q=0.3$ and $p=p_{max}=1.0$. Here $q<\alpha_{-}$, and crossover at $\alpha_{-}\approx 0.78$
The eigenvalues of $\chi(\alpha,q,p)$ are
$\Lambda_{\mathbb{I}}$ $=\frac{1}{4}+\frac{3}{4}\left[\frac{p\left(4+4\alpha-3\alpha p\right)-4}{4q+4\alpha q-3\alpha q^{2}-4}\right];$
$\Lambda_{i}$ $=\frac{1}{4}-\frac{1}{4}\left[\frac{p\left(4+4\alpha-3\alpha p\right)-4}{4q+4\alpha q-3\alpha q^{2}-4}\right],$ (15)
where $i=\{X,Y,Z\}$. In the case of no perturbations, $i.e.,$ $\alpha=0$ and $q=0$, the map is CP-divisible in the range $0\leq p\leq 1$. In this case, the eigenvalues become $\Lambda_{\mathbb{I}}=\left(1-\frac{3}{4}p\right)$, and $\Lambda_{i}=(\frac{p}{4})$. Furthermore, when $p=p_{max}=1.0$, the depolarizing is maximum, indicated by the intersecting points of the black lines in Fig. (1).
On the other hand, introducing a perturbation during the dynamics, initiates some form of memory into the system. These memory effects are controlled by the factor $\alpha$. The eigenvalues of the Choi matrix show crossover for $\alpha\neq 0$ and $q\neq 0$. The timelike $p$ varies from $q$ to $p_{max}=1$. The $p$ satisfying $04p+4\alpha p-3\alpha p^{2}-4=$ 0 gives $\alpha_{\pm}=\frac{2(1+\alpha\pm\sqrt{1-\alpha+\alpha^{2}})}{3\alpha}$, corresponding to points of crossover of eigenvalues. The $\alpha_{+}$ is outside the domain of $p$ and is not considered for study. For $q<\alpha_{-}$, eigenvalues crossover at $\alpha_{-}$. As depicted in Fig. (1), the point $p=\alpha_{-}$ represents maximal depolarizing. For $q>\alpha_{-}$, as illustrated in Fig. (2), $\Lambda_{i}$ become negative in the entire range of $p$ for the intermediate map indicating non-Markovianity. The point $q=\alpha_{-}$ represents a singularity, since $\Lambda_{I}$ and $\Lambda_{X,Y,Z}$ diverge for any $p\in(q,1]$. The negative eigenvalue $\Lambda_{X,Y,Z}$ suggests that the trace norm of Choi matrix, $||\chi(\alpha,q,p)||_{1}$ is greater than 1 in this interval. Since $\Phi(t,s)$ is trace-preserving, $||\chi(\alpha,q,p)||_{1}$ provides a witness of the NCP character of $\Phi(t,s)$ as [7]
$||\chi(\alpha,q,p)||_{1}$ $=||\left[\Phi(t,s)\otimes\mathbb{I}\right]\ket{\Psi}\bra{\Psi}||_{1}$
[…]
Integrating $\chi(\alpha,q,p)$ over the evolution time provides a measure of non-Markovianity, the RHP measure [7]. Moreover, NCP intermediate map relates to negative decay rates, suggesting a conceptually equivalent, but quantitatively different and possibly computationally easier method to indicate non-Markovianity, based on the integral of the normalized time dependent decay rate function in the canonical form of the master equation for the negative decay rate(s), the HCLA measure [19], used below.
Figure 2: The eigenvalues of for the Choi matrix (III.1). Sold line ($\lambda_{\mathbb{I}}$) and, dashed line ($\lambda_{X,Y,Z}$) , for $\alpha=0.7$ (non-Markovian). The intermediate $p$-range lies between $q=0.8$ and $p_{max}=1.0$. The whole range of $p$ corresponds to an NCP map, demonstrating the non-Markovianity of the channel. Here $q>\alpha_{-}$ as for $\alpha=0.7$ corresonding $\alpha_{-}\approx 0.78$.
III.2 The negativity of the decay rate
As stated in the preceding subsection, an NCP intermediate map is related to the observation of the negativity of the time-dependent decay rate $\gamma(p)$ of the master equation in the canonical form.
In some cases of interest, the open system dynamics may be written in terms of a time local master equation involving time-dependent functions as prefactors with otherwise Gorini-Kossakowski-Sudarshan-Lindblad (GKSL) form [33, 34]. Then, for some times during the system evolution, negative decay rates may show up, indicative of the non-Markovianiaty of the dynamics. The generalised Markovian dynamics appears when the master equation takes the quasi-GKSL-form [35]
$\frac{d\rho(p)}{dp}$ $=-i\left[H,\rho(p)\right]+\sum_{i}\gamma_{i}(p)\left(L_{i}(p)\rho(p)L_{i}^{\dagger}(p)\right.$
$\left.-\frac{1}{2}\left\{L_{i}^{\dagger}(p)L_{i}(p),\rho(p)\right\}\right),$ (17)
written in a canonical form [19], where the $L_{i}(p)$ are trace-less orthonormal operators. The decay rates $\gamma_{i}(p)\geq 0$ defines a CP-divisible dynamical map. It is natural to regard dynamical maps $\Phi(t,0)$ with master equations of type Eq. (III.2) for which $\gamma_{i}(p)<0$ for some $i$ and some time as candidates for non-Markovian quantum dynamics [35]. In these cases, the dynamical map is no longer CP-divisible.
Figure 3: (Color online) A plot of decoherence rates, $\gamma_{X,Y,Z}(p)$, as a function of $p$ for $\alpha=0.0$ (red dashed curve) and $\alpha=0.7$ (blue solid curve).
The Kraus representation Eq. (II) is a solution to the master equation describing the depolarizing map in the canonical form:
$\frac{d\rho}{dp}=\sum_{i=X,Y,Z}\gamma_{i}(p)\left[\sigma_{i}\rho(p)\sigma_{i}-\rho(p)\right],$ (18)
[…]
$\gamma(p)=-\frac{1}{G}\frac{dG}{dp}=\frac{4+(4-6p)\alpha}{4+3\alpha p^{2}-4p(1+\alpha)}.$ (19)
Figure (3) illustrates a plot of decay rates $\gamma_{X,Y,Z}(p)$ as a function of $p$ for $\alpha=0.0$ (red dashed curve) and $\alpha=0.7$ (blue solid curve). We observe that when the non-Markovian (perturbation) parameter, $\alpha$, is equal to zero, the decay rates become positive for the entire range of $p$ and then the dynamics is CP-divisible, as expected. Furthermore, when $\alpha=0.7$, the decay rates turn into negative at the singularity point, $i.e.,~{}p\approx 0.78$, indicating non-Markovian dynamics.
Using the canonical decay rates, $\gamma_{i}(p)$, and the master equation, Eq. (18), a function $f_{i}(p)$ is defined [19],
$f_{i}(p)=\{-\gamma_{i}(p),0\}.$ (20)
If $f_{i}(p)=0$, at any time, then the evolution is CP-divisible. The functions $f_{i}(p)$ can be used to quantify the non-Markovianity as
$\mathcal{N}_{HCLA}=\int_{p=q}^{p=p_{max}}-\gamma(p)dp,$ (21)
where $\gamma(p)$ is as in Eq. (19). It diverges because of the singularity of $\gamma(p)$ at $p\approx 0.8$ (see Fig. (3). One solution to this problem, following an idea proposed in [7, 20], is to replace $\gamma(p)$ by its normalized form
$\tilde{\gamma}(p)$ $=\frac{-\gamma(p)}{1-\gamma(p)}$
$=\frac{4+4\alpha-6\alpha p}{4p+4\alpha-2\alpha p-3\alpha p^{2}}.$ (22)
Hence,
[…]
=$=\Biggl{\{}\log\left(2p(-2+\alpha)-4\alpha+3p^{2}\alpha\right)$
$+\frac{6\alpha(tanh^{-1}\left(\frac{-2+\alpha+3p\alpha}{\sqrt{4-4\alpha+13\alpha^{2}}}\right))}{\sqrt{4-4\alpha+13\alpha^{2}}}\Biggr{\}}^{p=1}_{p=\alpha_{-}}.$ (23)
Figure (4) represents the values of $\mathcal{N}^{{}^{\prime}}_{HCLA}$ (blue, solid curve) as a function of non-Markovian (perturbation) parameter $\alpha$. The monotonic increase of this measure with $\alpha$ indicates non-Markovianity. These results are connected to the RHP measure, $\mathcal{N}_{RHP}$, since both coincide for the two level system [19].
Figure 4: (Color online) A plot of $BLP$ measure $\mathcal{N}_{BLP}$ (red, dashed line), and $HCLA$ measure $\mathcal{N}^{{}^{\prime}}_{HCLA}$ (blue, solid curve).
III.3 The distinguishability of quantum states
The BLP method observes the distinguishability between two initial states, $\rho^{S}_{1}$ and $\rho^{S}_{2}$ by calculating the $trace$ $distance$, $D$, between them. $D$ is a metric quantity on the space of physical states, satisfying $0\leq D\leq 1$, where $D=1$ if and only if $\rho^{S}_{1}$ and $\rho^{S}_{2}$ have orthogonal support, and characterizes a realistic upper bound for the distinguishability between the probability distributions resulting from measurements executed on the quantum states. Therefore, the trace distance can be given an interpretation as the distinguishability between two quantum states [5, 3]. If $\Phi(t,0)$ is CP map, then, it is “CP-divisible” and furthermore is a contraction for the trace distance, $i.e.,$ $D$ is non-increasing under CP-divisible maps,
$D\left[\Phi(t,0)(\rho^{1}_{S}),\Phi(t,0)(\rho^{2}_{S})\right]\leq D\left[\rho^{1}_{S},\rho^{2}_{S}\right],$ (24)
where in general,
$D\left[\rho^{2}_{S},\rho^{1}_{S}\right]=\frac{1}{2}||\rho^{2}_{S}-\rho^{1}_{S}||_{1}=\frac{1}{2}Tr\sqrt{(\rho^{2}_{S}-\rho^{1}_{S})^{2}}.$ (25)
The violation of Eq. (24) would be an indicator of non-Markovian behavior [5]. As an example, we calculate the trace distance between two different initial states, with orthogonal support, which evolve under depolarizing dynamics, Eq. (III.1). Let, $\rho^{1}_{S}(0)=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix},$ and $\rho^{2}_{S}(0)=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$. By applying the depolarizing channel, Eq. (III.1), on these initial states, the final states turn out to be
$\rho^{1}_{S}(t)$ $=\Phi(t,0)\left[\rho_{S}^{1}(0)\right]$
$=\frac{1}{2}\begin{pmatrix}\frac{1}{2}&\frac{p}{2}+\frac{\alpha p}{2}-\frac{3}{8}\alpha p^{2}-\frac{1}{2}\\ \frac{p}{2}+\frac{\alpha p}{2}-\frac{3}{8}\alpha p^{2}-\frac{1}{2}&\frac{1}{2}\end{pmatrix},$ (28)
$\rho^{2}_{S}(t)$ $=\Phi(t,0)\left(\rho_{s}^{2}(0)\right)]$
$=\frac{1}{2}\begin{pmatrix}\frac{1}{2}&-\frac{p}{2}-\frac{\alpha p}{2}+\frac{3}{8}\alpha p^{2}+\frac{1}{2}\\ -\frac{p}{2}-\frac{\alpha p}{2}+\frac{3}{8}\alpha p^{2}+\frac{1}{2}&\frac{1}{2}\end{pmatrix}.$ (31)
By substituting Eq. (III.3) into Eq. (25), we obtain
$D\left[\rho^{2}_{S},\rho^{1}_{S}\right]$ $=\frac{1}{2}Tr\sqrt{(\rho^{2}_{S}-\rho^{1}_{S})^{2}}$
$=\frac{1}{4}|4+3\alpha p^{2}-4p(\alpha+1)|.$ (32)
Figure 5: (Colour online) Log plot of trace distance $D$ for $\alpha=0.0$ (solid red curve), $\alpha=0.7$ (blue dashed curve), and $\alpha=0.9$ (orange dashed-dotted curve), as a function of $p$. Note that, $\alpha=0.0$ describes the Markovian dynamics since $D$ decreases monotonically for all values of $p$.
The log plot of trace distance $D$ as a function of $p$ with $\alpha=0$ (solid red curve), $\alpha=0.7$ (blue dashed curve), and $\alpha=0.9$ (orange dashed-dotted curve), under the considered non-Markovian depolarizing noise is depicted in Fig. (5). Note that larger $\alpha$ shows a larger enhancement region, suggesting larger non-Markovianity in the sense of BLP [5]. The BLP measure of non-Markovianity, $\mathcal{N}_{BLP}$, is given by
[…]
If $\frac{dD}{dp}>0$ for some $p$, then the dynamics is non-Markovian [7], and the quantification of this non-Martkovianity can be done by calculating $\mathcal{N}_{BLP}$
[…]
$\mathcal{N}_{BLP}$ is illustrated in Fig. (4) (red, dashed line) and shows an agreement with the quantification of non-Markovianity according to the normalised $\mathcal{N}^{{}^{\prime}}_{HCLA}$.
III.4 Using quantum memory witness
The intermediate dynamical map $\Phi(t,s)$ and the corresponding Choi matrix $\chi(\alpha,q,p)$ are useful for the observation and certification of non-Markovian memory effects. The breaking of CP divisibility can be captured using the Choi matrix, which is Hermitian, allowing one to write a spectral decomposition as $\chi(\alpha,q,p)=\sum_{i}\lambda_{i}P_{i}$. The $P_{i}$ are orthogonal projections derived using corresponding eigenvectors, and the $\lambda_{i}$ are eigenvalues associated with $\chi(\alpha,q,p)$. The occurrence of negative $\lambda_{i}$ points to non-Markovianity. Following this, the Bell states (|$(\ket{\phi^{+}},\ket{\psi^{\pm}})$ are observed to serve as witness operators for depolarizing channel [37] giving
$tr(\ket{\phi^{+}}\bra{\phi^{+}}\chi(\alpha,q,p))=\Lambda_{I},$
$tr(\ket{\psi^{\pm}}\bra{\psi^{\pm}}\chi(\alpha,q,p))=\Lambda_{i=x,y,z}.$ (35)
One of the eigenvalues ($\Lambda_{I},\Lambda_{i=x,y,z}$) becomes negative for $q=0.3$ and $\alpha>0.7$, indicating that the intermediate evolution is not CP divisible in the interval [$q,p$] and hence non-Markovian. The map admits information backflow in this interval, and usual tools lack the ability to distinguish this information as classical or quantum in nature. We use [21] to characterise the quantum nature of information backflow in the intermediate depolarizing map. The existence of the Hermitian witness operator, in the form of the Bell states, then allows the establishment of the variable $X_{\chi(\alpha,q,p)}$ corresponding to $\chi(\alpha,q,p)$. This is used to ascertain the quantum nature of the information backflow, characterized by non-monotinicity of $X_{\chi(\alpha,q,p)}$, by satisfying the following condition [21]
$X_{\chi(\alpha,q,p)}=|\vec{s}|+||T||_{1}>1,$ (36)
where $\vec{s}$ consists of $s_{i}=tr(\chi(\alpha,q,p)(I\otimes\sigma_{i}))$, and $T$ is a 3$\times$3 matrix consisting of elements $T_{ij}=tr(\chi(\alpha,q,p)(\sigma_{i}\otimes\sigma_{j}))$. The expression for $X_{\chi(\alpha,q,p)}$ is
$X_{\chi(\alpha,q,p)}=3\frac{\lvert p\left(4+4\alpha-3\alpha p\right)-4\rvert}{\lvert 4q+4\alpha q-3\alpha q^{2}-4\rvert}.$ (37)
It quantifies the memory effects, and any non-monotonic increase in $X_{\chi(\alpha,q,p)}$ can be ascribed to quantum information backflow in interval [$q,p$]. The Fig. (6) depicts the variation of quantum memory effects using $X_{\chi(\alpha,q,p)}$. For $q=0.3$ and $\alpha=0.8$, Fig. (6) depicts a non-monotonic increase, implying information backflow which is quantum in nature and not classical [21]. Further non-monotonic variation of $X_{\chi(\alpha,q,p)}$ increases as $\alpha$ increases, indicating enhanced quantum information backflow, hence non-Markovianity. The increase in value of $X_{\chi(\alpha,q,p)}$ is due to the system recovering information, particularly quantum information lost to the environment.
Figure 6: (Color online) The variation of $X_{\chi(\alpha,q,p))}$, Eq. (37), for $q=0.3,\alpha=0.$ (red solid line), $q=0.3,\alpha=0.7$ (blue dashed line), $q=0.3,\alpha=0.8$ (orange dashed-dot line),for $q=0.3,\alpha=0.9$ (purple dashed line), and $q=0.3,\alpha=1$ (green large-dashed line).
IV Non-Markovianity via Geometrical visualization techniques
In this section, we demonstrate the geometrical observation of the non-Markovian behavior of the depolarizing map through witnessing and visualizing the varying of the rate of the volume of physical states of the system affected by the depolarizing dynamics. Furthermore, we visualize the divisibility of the depolarizing map and the conditions for the non-Markovian dynamics of the depolarizing channel.
IV.1 Non-Markovianity through volume change of physical states of the system
Figure 7: Time evolution of the $|M(p)|$ for a generic Markovian (left graph) and non- Markovian dynamics (right graph). In the Markovian case, the Bloch sphere shrinks along $x$, $y$, and $z$ axis and so the volume decreases monotonically for all values of $p$. For the non-Markovian case, the volume grows suddenly at $p\approx 0.7$ indicating the non-Markovianity beyond this point.
Rate of change in the volume of accessible states to the system is a tool in [22], enabling a geometrical visualization of the dynamical effects of non-Markovianity. Unlike other methods, see, for example, [5], this method need not to be optimized over the initial states. If the quantum evolution is Markovian, then this implies that the domain’s volume of the dynamical map reduces monotonically, as indicated in the left side of Fig. (6). In contrast, for a non-Markovian dynamical map, there could be an increase in the domain’s volume for some time interval during the dynamics.
The density matrix of a qubit can be written using operators $G=\frac{1}{\sqrt{2}}\{I_{0},\sigma_{x},\sigma_{y},\sigma_{z}\}$ and generalized Bloch vector ($\vec{r}$) [3, 25] as
$\rho=\sum_{i=0}^{3}r_{i}G_{i}.$ (38)
Here, $r_{i}=tr(G_{i}\rho)$ such that $\vec{r}=(\frac{1}{\sqrt{2}},r_{i=1,2,3})$. If a map $\Phi$ acts on a single qubit then $\Phi:\rho\rightarrow\rho^{\prime}$, where $\rho^{\prime}$ must be expressible in terms of a new Bloch vector $\vec{r}^{\prime}$ associated to $\rho^{\prime}$. The action of a unital quantum map $\Phi$ on a qubit state can be represented by mapping the Bloch vector according to $\vec{r}\rightarrow\vec{r}^{\prime}=M\vec{r}$. The equation, $M_{ij}=tr[G_{i}\Phi[G_{j}]]$, gives the respective matrix elements. This is an affine transformation for the Bloch vector, for the respective $M$, as
$M=\begin{pmatrix}1&0&0&0\\ 0&\lambda_{1}(p)&0&0\\ 0&0&\lambda_{2}(p)&0\\ 0&0&0&\lambda_{3}(p)\end{pmatrix},$ (39)
where, $\lambda_{1}(p)=\lambda_{2}(p)=\lambda_{3}(p)=\frac{3}{4}\alpha p^{2}-\alpha p-p+1$. The matrix $M$ can be written as
$M=\begin{pmatrix}1&\textbf{0}\\ \textbf{0}&\textbf{B}\end{pmatrix}.$ (40)
Here, B is the matrix satisfying condition $|\mathbf{B}|=|M|$. The absolute value of determinant of matrix $M$, $|M|$, can be associated with volume of accessible states [38]. It gives the reduction factor for the volume of accessible states, given by the measure of the set of evolved Bloch vectors, with respect to its value at $p=0$. We plot $|M(p)|$ as a function of $p$ for the evolution of the Bloch vectors under the action of the general depolarizing noisy map in Fig. (7) for $\alpha=0$ (no perturbation), left and $\alpha=0.8$ (with perturbation), right. When $\alpha=0$, the evolution of the norm of the map demonstrates Markovian dynamics, as can be discerned from the monotonic decrease in $|M(p)|$. For $\alpha=0.8$, increase in the value of $|M(p)|$ is observed, for $p$ ranging from $0.7$ to $1.0$,indicating non-Markovianity. In spherical co-ordinates, it is then straightforward to check that any positive trace-preserving map described by Eq. (37) induces the change
$\frac{dV}{dp}=||\mathbf{B}||\cdot\frac{dV}{dp}|_{p=0},$ (41)
where $||\mathbf{B}||$ decreases monotonically for any positive, linear, and trace preserving map [39, 22]. The values of $||\textbf{B}||$ as well as $||M(p)||$ could be used to identify non-Markovianity, as a similar trend like $|M(p)|$ will be observed for them. The $\frac{d||M(p)||}{dp}>0$ for some $p$, is then indicative of non-Markovian dynamics. Using this, the quantification of non-Markovianity for the depolarizing channels is
[…]
It is worth mentioning that $\mathcal{N}_{BLP}$ and $\mathcal{N}_{v}$ measures show similarity.
IV.2 Non-Markovianity through visualizing the trajectory traced by map in the parameter space
The diagonalized form of the matrix elements of the intermediate map $\Phi(t,s)$ can be defined by three real parameters $\lambda_{1}(p),\lambda_{2}(p)$, and $\lambda_{3}(p)$, as shown in Eqs. (11) and (12). An arbitrary Pauli map is defined as [40, 23]
$\Phi(\rho)$ $=\frac{1}{2}\left(tr(\rho)\mathbb{I}+\sum_{i=1}^{3}\lambda_{i}(p)tr(\sigma_{i}\rho)\sigma_{i}\right)$
$=\begin{pmatrix}1&0&0&0\\ 0&\lambda_{1}(p)&0&0\\ 0&0&\lambda_{2}(p)&0\\ 0&0&0&\lambda_{3}(p)\end{pmatrix},$ (47)
where $\sigma_{1}$, $\sigma_{2}$, $\sigma_{3}$ is a regular set of Pauli operators. Following [40, 23] the map $\Phi(\rho)$ is positive if $-1\leq\lambda_{i}\leq 1$ (cube in the parameter space) and the map $\Phi(\rho)$ is CP if $1\pm\lambda_{3}(p)\geq|\lambda_{1}(p)\pm\lambda_{2}(p)|$. The smooth process $\Phi(\rho)$ can then be determined by a continuous trajectory ${\bf{\lambda}}(p)$ in the parameter space. Such a trajectory provides a pictorial representation of the dynamical map in $\mathbb{R}^{3}$. By analysing the process trajectory in the parameter space, its (non-)Markovian properties are revealed. In [23], the dynamical maps were visualised via paths in the parameter space and their divisibility was investigated. The dynamics looses CP-divisibility under infinitesimal perturbations. Any qubit Pauli map $\Phi(\rho)$ between finite dimensional spaces can be described by three real parameters $\left(\lambda_{1}(p),\lambda_{2}(p),\lambda_{3}(p)\right)$ [19, 41, 42, 43]. Investigating the process trajectory in the parameter space brings out its divisibility properties. Our goal here is to detect the non-Markovianity of the depolarizing map in terms of its trajectory in the parameter space [13, 44, 23].
For any linear map, $\Phi(t,0)$, the trajectory $\{\lambda_{1},\lambda_{2},\lambda_{3}\}$ can be an arbitrary smooth curve inside the tetrahedron $1\pm\lambda_{3}\geq|\lambda_{1}\pm\lambda_{2}|$, Fig. (7). The curve starts initially, at $p=0$, from the corner $(\lambda_{1}(0)=1,\lambda_{2}(0)=1,\lambda_{3}(0)=1)$. If the traced trajectory curve stays in the tetrahedral part of the positive octant made from the cube shown in Fig. 7, the corresponding map $\phi_{t}$ is said to be CP. The trajectory of the curve residing in this positive octant pertains to some map $\phi_{t}$ with positive decoherence rates. The curve traversing out of this octant is indicative of a negative decoherence rate and non-CP behaviour [23].
Figure 8: Geometrical presentation of the map where the curve corresponds to CP maps since it is located inside the tetrahedron. The vector A(p) is drawn from corner (1,1,1) and is pointing inside the tetrahedron, indicating the ”CP-divisibility.
((a))
((b))
Figure 9: Tragectory-based visualization of dynamics (a) for $\alpha=0$; (b) for $\alpha=0.7$.
To discover direction during the dynamical progress in the parameter space, we use the following formula [23]
$\dot{\Phi}o\Phi^{-1}\left(\rho\right)=\frac{1}{2}\sum_{i=1}^{3}\frac{\dot{\lambda_{i}}(p)}{\lambda_{i}(p)}tr(\rho)\sigma_{i}$ (48)
Accordingly, we define a vector $\bf{A}(p)$ as
${\bf{A}}(p)=\left(\frac{\dot{\lambda}_{1}(p)}{\lambda_{1}(p)},\frac{\dot{\lambda}_{2}(p)}{\lambda_{2}(p)},\frac{\dot{\lambda}_{3}(p)}{\lambda_{3}(p)}\right),$ (49)
where, $\dot{\lambda}_{i}(p)\equiv\frac{d{\lambda}_{i}(p)}{dp}$. The vector $\bf{A}(p)$ can be established at any value $p$, highlighting the non-Markovian property of the dynamics. The dynamic is CP-divisible if and only if the vector $\bf{A}(p)$ drawn from the corner $(1,1,1)$ of the parameter space points inside the tetrahedron in Fig. (8). The scalar products of $\bf{A}(p)$ with vectors $(-1,1,1)$, $(1,-1,1)$, and $(1,1,-1)$ are all non-positive
$-A_{1}(p)+A_{2}(p)+A_{3}(p)\leq 0,$ (50a)
$A_{1}(p)-A_{2}(p)+A_{3}(p)\leq 0,$ (50b)
$A_{1}(p)+A_{2}(p)-A_{3}(p)\leq 0.$ (50c)
If a vector $\bf{A}(p)$ fails to satisfy Eq. ((50a)-(50c)), it indicates the non-Markovian CP-indivisible nature of the dynamics. For a general depolarizing map, its trajectory is given by $\lambda(p)=(\lambda_{1}(p),\lambda_{2}(p),\lambda_{3}(p))=\frac{3}{4}\alpha p^{2}-\alpha p-p+1$. In the case of Markovian dynamics, where $\alpha=0$ (no perturbation), the expressions become $A_{1}(p)=A_{2}(p)=A_{3}(p)=\left(1-p\right)$, and $\dot{\lambda}_{1}(p)=\dot{\lambda}_{2}(p)=\dot{\lambda}_{3}(p)=-1$. This results in $A_{1}(p)=A_{2}(p)=A_{3}(p)=\frac{-1}{1-p}<0$, for $p\in[0,1]$. The trajectory curve, depicted in Fig. (8a), is inside the positive octant; the curve traverses a path from $(1,1,1)$ to $(0,0,0)$. Consequently, the map $\Phi(t,0)$ satisfies Eq. (50a)-(50c), indicating that the dynamics is CP-divisible. Conversely, when $\alpha=0.7$ (with perturbation), the curve is modified to visualise the absolute value of eigenvalues. It is then observed that the curve traverses a reverse path for $p>\alpha_{-}(\alpha=0.7)$. This shift is shown in the plot given in Fig. (8b). The expressions for $A_{1}(p)$, $A_{2}(p)$, and $A_{3}(p)$ become $\left(\frac{42p-68}{21p^{2}-68p+40}\right)$. The inequalities in Eq. (42)-(48) are violated for the intermediate map when $0.8\leq p\leq 1$. Hence, $\Phi(t,0)$ is non-Markovian in this interval, in consistency with our findings in the previous section.
V Conclusion
The notion of non-Markovianity in completely positive depolarizing noise is explored using various dynamical and geometrical techniques. The introduced non-Markovianity parameter ($\alpha$) of the depolarizing map, is the cause of perturbations and non-Markovianity. It contributes to breaking the CP divisibility of the intermediate map, assuring non-Markovianity. The non-Markovian behaviour of the depolarizing map could be identified by using appropriate witness operators. The intermediate evolution admits information backflow, showing non-monotonic variation in quantum memory, which could be justified using witness $X_{\chi(\alpha,q,p)}$. The witnessed non-Markovianity is quantified using two techniques: (a) BLP measure based on the distinguishability of states, (b) HCLA measure based on negativity of decoherence rates. We further study geometric visualisation techniques to gain a comprehensive understanding of non-Markovianity. It is observed that non-Markovianity contributes to enhancing the accessible volume of states during evolution. Furthermore, the trajectory traced during evolution in parameter space presents the loss of the CP divisibility under perturbations caused by non-zero $\alpha$.